OpenStudy (anonymous):

How can you write the expression with a rationalized denominator? (sqrt3-sqrt6) divided by (sqrt3+sqrt6)

jimthompson5910 (jim_thompson5910):

multiply top and bottom by \[\Large \sqrt{3}-\sqrt{6}\]

OpenStudy (anonymous):

so i got ... 9 – 2 sqrt18

OpenStudy (anonymous):

is that right ?

jimthompson5910 (jim_thompson5910):

\[\Large (\sqrt{3}-\sqrt{6})(\sqrt{3}-\sqrt{6})\] \[\Large (\sqrt{3})^2-2*\sqrt{6}*\sqrt{3}+(\sqrt{6})^2\] \[\Large 3-2*\sqrt{6*3}+6\] \[\Large 3-2*\sqrt{18}+6\] \[\Large 9-2*\sqrt{18}\] Thats in the numerator. In the denominator you have 3 - 6 = -3 So the fraction becomes \[\Large \frac{9-2*\sqrt{18}}{-3}\] So the final answer would be \[\Large -\frac{9-2*\sqrt{18}}{3}\] which can be written as \[\Large \frac{-9+2*\sqrt{18}}{3}\]

OpenStudy (anonymous):

thats not an option.. all there is.. 9 – 2 sqrt18 and -3 -2*sqrt18 divided by 9

OpenStudy (anonymous):

and -1-2*sqrt18 divided by 3

jimthompson5910 (jim_thompson5910):

\[\Large \frac{-9+2*\sqrt{18}}{3}\] \[\Large \frac{-9}{3}+\frac{2*\sqrt{18}}{3}\] \[\Large -3+\frac{2*\sqrt{18}}{3}\]

OpenStudy (anonymous):

thats still not my choices

jimthompson5910 (jim_thompson5910):

\[\Large 9-2*\sqrt{18}\] only works if it's divided by -3

OpenStudy (anonymous):

so would that be it ???

jimthompson5910 (jim_thompson5910):

well there must be a typo somewhere